### v1

# Steering Elongate Multi-legged Robots By Modulating Body Undulation Waves

Abstract
Centipedes exhibit great maneuverability in diverse environments due to their many legs and body-driven control. By leveraging similar morphologies, their robotic counterparts also demonstrate effective terrestrial locomotion. However, the success of these multi-legged robots is largely limited to forward locomotion; steering is substantially less studied, in part due to the challenges in coordinating their many body joints. Furthermore, steering behavior is complex and can include different combinations of desired rotational/translational displacement. In this paper, we explore steering strategies in multi-legged robots based on tools derived from geometric mechanics (GM). We characterize the steering motion in the plane by the rotation angle, the steering radius, and the heading direction angle. We identify an effective turning strategy by superimposing two traveling waves in the lateral body undulation and further explore variations of the “turning wave” to enable a broad spectrum of steering behaviors. By combining an amplitude modulation and a phase modulation, we develop a control strategy for steering behaviors that enables steering with a range of rotation angles (from $0^{\circ}$ to $20^{\circ}$) and steering radius (from 0.28 to 0.38 body length) while keeping the heading direction angle close to 0. Lastly, we test our control framework on an elongate multi-legged robot model to verify the effectiveness of our proposed strategy. Our work demonstrates the generality of the two-wave template for effective steering of multi-legged elongate robots.
I Introduction
Figure 1: (A.1) Ackerman steering strategy in a 4 wheeled vehicle utlizing changes to wheel orientation. (A.2) Differential drive steering strategy utilizing rotational velocity imbalance to induce turns on a wheeled vehicle. (B) Centipede robot steering strategy powered by body shape changes similar to Ackerman steering. (C) Horizontal body undulation joint index and angle definition
Multi-legged robots can traverse rugged landscapes without the need for extensive sensing due to their many legs, making them appealing for various sectors such as search-and-rescue, terrestrial exploration, and agriculture. legged systems such as quadrupeds and bipeds, where effective locomotion relies on the precise control of each leg, multi-legged robots can continue to successfully locomote after individual leg failure or missed contacts because of their morphological redundancy [1, 2].
Despite this robustness, whole-body coordination in multi-legged robots (e.g., body bending, leg stepping, and leg rotation) presents a high degree-of-freedom (DoF) problem that increases in complexity with the number of legs. Schemes have been developed to coordinate the many degrees of freedom for effective locomotion. One such scheme involves the propagation of a travelling wave down the body of the centipede to control the shape of the robot [2]. By coupling leg stepping to body shape, locomotion can be controlled through shape changes to the body, thus controlling the high DoF system via a lower order representation that is not limited by number of segments [3].
However, the success of controlling a multi-legged robot has been limited to generating straight, forward motion. There exists a significant research gap in producing more complex paths for steering centipede robots, which severely limits the applicability of these multi-legged systems. Having the ability to follow arcs (i.e., steer) is essential for effective autonomous navigation in obstacle-rich environments where the robot has to detour around objects. We are inspired by effective steering strategies in wheeled systems and aim to develop effective steering strategies for elongate multi-legged robots (Fig. 1B). Traditional approaches to steering wheeled systems involve geometric changes to the physical shape of the wheeled system (Ackerman, Fig. 1A.1) or differential drive using the imbalance of angular velocities to create arcs (Differential drive, Fig. 1A.2).
The differential drive approach is widely adapted in legged locomotion, especially for few-legged robots such as quadrupeds or bipeds. Previous work has shown that careful foot placement planning can enable quadrupedal turning motions ([4]). Cho et al., [5] introduced the footfall planning objective function, where speed, stability, translation direction, and turning were all included. Palmer et al., [6] designed a turning fuzzy controller by placing the fore legs to the outside of the turn and the hind legs to the inside.
However, the differential drive approach is not readily applicable to elongate multi-legged robots partially because of their relatively simple legs (e.g., mechanical constraints on leg actuation) and their redundant ground reaction force (the differential torque from each leg pair can cancel with each other). Instead, we propose a steering scheme similar to Ackerman steering by altering the body undulation. The challenges lie in coordinating not only many but also different types of joint movements (lateral/vertical body bending, leg stepping/protraction, etc.).
In this paper, we will use geometric mechanics (GM) to resolve these challenges. GM has been developed as a general scheme to understand the principles behind seemingly complicated locomotion behaviors [7][8, 9, 10, 11, 12]. Specifically, the motion of a locomoting system is separated into a shape space (e.g., the internal joint angle space 1.C) and a position space (position and orientation of a locomotor in the world frame). By establishing a mapping between velocities in shape and position spaces, geometric mechanics offers tools that allow us to visually analyze, design, and optimize gaits.
Using geometric mechanics, prior work has identified separate body-driven turning strategies for quadrupedal robots [13] and elongate limbless robots [14, 15]. Centipede robots contain features present in both these systems, and thus, we posit that a body-driven turning strategy can be applied to multi-legged devices using geometric mechanics. We combine the two turning strategies (quadruped-like and limbless-like) and propose a simple control that is capable of generating a wide variety of circular arcs ( $0^{\circ}$ to $20^{\circ}$ per cycle). Our results significantly enrich the library of locomotion strategies for elongate multi-legged robots and provide effective tools to navigate around obstacles, paving the way towards all-terrain, highly-capable, elongate multi-legged robots.
II Geometric mechanics
II-A Two-basis function turning
We first create a low-dimensional space for multi-legged locomotion. As documented in prior work [2, 16], we assume that the lateral body undulation wave can be prescribed by:
$\alpha(i,t)$ $=w_{1}(t)S_{1}(i)+w_{2}(t)S_{2}(i)$
$S_{1}(i)$ $=\sin{(2\pi\frac{i}{N})}$
$S_{2}(i)$ $=\cos{(2\pi\frac{i}{N})}$ (1)
where $\alpha(i,t)$ denotes the yaw angle of $i$-th joint at time $t$ (Fig. 1C); $S_{1}(i)$ and $S_{2}(i)$ are shape basis functions; $N$ is the number of leg pairs; $w_{1}(t)$ and $w_{2}(t)$ are the time series of weights for the corresponding shape basis functions. Here, we define the shape variable as $\boldsymbol{w}(t)=[w_{1}(t),w_{2}(t)]$, which then uniquely characterizes the body movement. As discussed in prior work [2, 16], the contact pattern of legs and the leg shoulder angles can also be prescribed by the shape variable $\boldsymbol{w}$ (detailed prescription equations can be found in [2]). In Fig. 2.A, we illustrated the shape space for a 10-legged robot.
Perturbation in the shape variable can result in net displacement. The net translation in the plane, SE(2), can be characterized by $\Delta x,\ \Delta y,\ \Delta\theta$ in forward, lateral, and rotational directions, respectively. To better characterize turning/steering in SE(2), we use the following notation to describe the net translation in the plane: $r$, $\Delta\theta$, and $\gamma$, which characterize the steering curvature, rotational translation, and changes in the heading directions, respectively (Fig. 4). Notably, for small perturbations in the shape variables (thus, small displacements), we have: $r\Delta\theta=|[\Delta x,\Delta y]|$, and $\tan(\gamma)=\Delta y/\Delta x$.
We define the body velocity, $\boldsymbol{\xi}=[\xi_{x},\xi_{y},\xi_{\theta}]$, as the overall locomotor velocity in the forward, lateral, and rotational directions [17]. Specifically, we have: $\boldsymbol{\xi}=\lim_{t\rightarrow 0}\frac{[\Delta x,\ \Delta y,\ \Delta\theta]}{dt}$. We can then numerically calculate the body velocity from shape variables ($\boldsymbol{w}$) and the shape velocity ($\dot{\boldsymbol{w}}$) via net ground reaction forces (GRF) analysis. Here, we model the ground reaction force by Coulomb friction. From geometry, the GRF at each foot can be uniquely expressed as a function of shape variable ($\boldsymbol{w}$), shape velocity ($\dot{\boldsymbol{w}}$), and body velocity ($\boldsymbol{\xi}$). Assuming quasi-static motion, we consider the total net force applied to the system is zero at any instant in time:
$\boldsymbol{F}=\sum_{i\in I_{\boldsymbol{w}}}{\left[\boldsymbol{F}^{i}_{\parallel}\left(\boldsymbol{\xi},\boldsymbol{w},\dot{\boldsymbol{w}}\right)+\boldsymbol{F}^{i}_{\perp}\left(\boldsymbol{\xi},\boldsymbol{w},\dot{\boldsymbol{w}}\right)\right]}=0,$ (2)
where $I_{\boldsymbol{w}}$ is the collection of all stance-phase legs, determined by the shape variable $\boldsymbol{w}$ [2]. At a given body shape $\boldsymbol{w}$, Eq.(2) connects the shape velocity $\dot{\boldsymbol{w}}$ to the body velocity $\boldsymbol{\xi}$. Therefore, by the implicit function theorem and the linearization process, we can numerically derive the fundamental equation of motion:
$\boldsymbol{\xi}~{}\approx\boldsymbol{A}(\boldsymbol{w})\dot{\boldsymbol{w}}=\begin{bmatrix}\boldsymbol{A}^{x}(\boldsymbol{w})\\ \boldsymbol{A}^{y}(\boldsymbol{w})\\ \boldsymbol{A}^{\theta}(\boldsymbol{w})\end{bmatrix}\dot{\boldsymbol{w}},$ (3)
where $\boldsymbol{A}$ is the local connection matrix, $\boldsymbol{A}^{x},\boldsymbol{A}^{y},\boldsymbol{A}^{\theta}$ are the three rows of the local connections, respectively. Each row of the local connection matrix can be regarded as a vector field over the shape space, called the connection vector field. In this way, the body velocities in the forward, lateral, and rotational directions are computed as the dot product of connection vector fields and the shape velocity $\dot{\boldsymbol{w}}$.
Consider a gait as a periodic pattern of self-deformation: $\{\partial\phi\ :[w_{1}(t),w_{2}(t)],t\in(0,T]\}$, where $T$ is the temporal period. The displacement along the gait path $\partial\phi$ over a cycle can be approximated to the first order by:
$\begin{pmatrix}\Delta x\\ \Delta y\\ \Delta\theta\end{pmatrix}=\int_{\partial\phi}{\begin{bmatrix}\boldsymbol{A}^{x}(\boldsymbol{w})\\ \boldsymbol{A}^{y}(\boldsymbol{w})\\ \boldsymbol{A}^{\theta}(\boldsymbol{w})\end{bmatrix}\mathrm{d}\boldsymbol{w}}.$ (4)
Of our particular interest in turning, we have $\Delta\theta=\int_{\partial\phi}\boldsymbol{A}^{x}(\boldsymbol{w})\mathrm{d}\boldsymbol{w}$. According to Stokes’ Theorem, the line integral along a closed curve $\partial\phi$ is equal to the surface integral of the curl of $\boldsymbol{A}^{\theta}(\boldsymbol{w})$ over the surface enclosed by $\partial\phi$:
$\Delta\theta=\int_{\partial\phi}{\boldsymbol{A}^{\theta}(\boldsymbol{w})\mathrm{d}\boldsymbol{w}}=\iint_{\phi}{\nabla\times\boldsymbol{A}^{\theta}(\boldsymbol{w})\mathrm{d}w_{1}\mathrm{d}w_{2}},$ (5)
where $\phi$ denotes the surface enclosed by $\partial\phi$. The curl of the connection vector field, ×$\nabla\times\boldsymbol{A}^{\theta}(\boldsymbol{w})$, is referred to as the height function [18]. With the above derivation, the gait design problem is simplified to drawing a closed path in the shape space. Net displacement over a period can be approximated by the integral of the surface enclosed by the gait path. Hence, we are able to visually identify the optimal gait leading to the largest turning by finding the path with the maximum surface integral.
We illustrate the height function and an example gait path in Fig.2B. From the structures of height functions, we notice that (1) the positive and negative volumes are distributed symmetrically about the y-axis ($w_{1}=0$), and (2) turning can be induced if we introduce an offset to the center of the gait path (solid path in Fig. 2B). As documented in [3], adding an offset to the gait path in the shape space is an effective turning strategy for limbless locomotors such as nematode worms and sidewinders. However, we also noticed that the magnitude of the rotational height function is small, leading to an ineffective turning strategy. To verify our observation, we test the center-shifted path (the solid path in Fig. 2B) in simulation, and we note that the net turning over a cycle is less than $10^{\circ}$.
II-B Three-basis function turning
Figure 2: (A) The Euclidean shape space for a multi-legged robot representing a two-wave template. Robot body postures over the shape space is presented. (B) The rotational height functions using the two-wave template. Turning can be modulated by offsetting a gait path from the center (enclosing non-zero surface integral over the height function). Paths in shape space prescribe a series of robot shapes and coupled contact patterns that define a gait. (C) The simulation of the gait leading to net heading angle change less then $10^{\circ}$
Inspired by the nematode turning documentation [14, 15], we hypothesize that adding a third basis function can substantially improve the turning performance. Specifically, we introduce the third basis function:
$\alpha(i,t)$ $=w_{1}(t)S_{1}(i)+w_{2}(t)S_{2}(i)+w_{3}(t)S_{3}{i}$
$S_{3}(i)$ $=\cos(2\pi k\frac{i-N/2}{N})$ (6)
where $k$ is the spatial frequency of the third wave. For simplicity, we first consider $k=0$ in this section. The third basis function introduces a 3-dimensional shape space, which can present a substantial challenge for directly applying geometric mechanics [19, 20]. To simplify our analysis, we chose a fixed prescription on the trajectory in $w_{1}$-$w_{2}$ space:
$w_{1}$ $=\pi/6\cos(\phi)$ (7)
$w_{2}$ $=\pi/6\sin(\phi),$ (8)
where $\phi$ denotes the phase. Thus, we construct a new shape space $\boldsymbol{C}=[\phi,\ w_{3}]$. Notably, $\boldsymbol{C}$ has a cylindrical structure because $\phi$ is periodic (i.e., $\phi=0$ is equivalent to $\phi=2\pi$). We illustrate the new shape space in Fig. 3. Following similar steps as descried in Sec. II.A, we can obtain the height function in the new shape space (Fig. 3B). Notably, with the new height function, there is no apparent area enclosed for a given gait (e.g., the solid and dashed line in Fig. 3B), and we cannot readily perform the surface integral. Instead, prior works show that  [21, 22] the surface integral can be calculated as the area enclosed by the area underneath the curve.
We thus identify two effective turning gaits, represented by the solid and dash curves in Fig. 3B:
$w_{3}$ $=A_{3}\sin{(\phi)}\ \ \ or,$ (9)
$w_{3}$ $=A_{3}\sin{(\phi+\pi)}$ (10)
We verify the effectiveness using numerical simulation in Fig. 3C. Both gaits can lead to net rotation over $20^{\circ}$ per cycle.
We choose one of the effective turning gaits (Eq. 9) for further analysis. Specifically, we consider a modulation on the amplitude ($A_{3}$). We hypothesize that the modulation on this amplitude can enable a control over the net rotation $\Delta\theta$ and turning curvature $r$. To test our hypothesis, we performed the numerical simulation to estimate the net rotation $\Delta\theta$ and turning curvature $r$ as a function of $A_{3}$ (numerical simulation details can be found in [14, 15, 3]). As a result, we notice an almost linear relationship (Fig. 5.A, red curves) between the amplitude ($A_{3}$) and the net turning ($\Delta\theta$) or steering curvature ($1/r$).
Figure 3: (A) The cylindrical shape space for a multi-legged robot representing a three-wave template. Y-axis denote the amplitude of the third wave. The x-axis denotes the phase in the first-two waves. (B) The corresponding rotational height functions. We identified two effective steering gaits illustrated in solid and dash curves. The steering gaits can be prescribed by a sinusoidal function with an constant offset. (C) simulation of the two steering gaits.
III Experimental Setup
Figure 4: Centipede robot with 5 leg pairs used to test steering control scheme. (A) An illustration of a 10 -legged robot steering over two cycles. The red curve denotes the trajectory of the last leg on the left. Steering parameters are labeled: steering radius, $R$, heading angle, $\theta$, and orientation change angle from the tangent, $\theta$. (B) Markers are placed on the head and last module of the robot. The yellow curve denotes the marker trajectory over 3 cycles. The red curves are the fitted circle to estimate steering parameters. For clarity, the other marker trajectory was not displayed but was used in fitting the curve.
To test the possible turning behavior, we built a five segment centipede robophysical model with the ability to control horizontal and vertical body joint angles, leg rotation angles, and leg stepping contacts. Additionally, it has an on-board computer (Raspberry Pi 4), head module with a depth camera (Intel© RealSense) , and inertial measurement units on each link. Horizontal and vertical body undulation are controlled via Dynamixel 2XL430 servos while leg rotation and stepping motions use Dynamixel XL430 servos. Mechanical gears increase the torque of the body servos while rigid linkages control the stepping motion. There are ball casters on the leg contact to allow for free rotation on the carpet surface and to prevent sinking.
We define steering paths as circular arcs of fixed radius, $r$, where the heading angle change , $\theta$, describes the amount of the arc that is traversed and $\gamma$ describes the rotation of the robot relative to the tangent of the arc. A planar steering path can be fully defined by these three parameters for each robot cycle ($t=T$) as shown in Fig 4A.
We measured steering trajectories on the robot by tracking markers via an overhead camera (Fig 4B). To reduce noise and variability from the shape of the robot, the average position of two markers was used to capture the bulk motion of the robot, and the vector from tail to head was used to define the pose of the robot body. A circle is fit to the trajectory of the robot after completing 3 cycles to obtain $r$ while $\theta$ is obtained by taking the angle between the initial and final pose of the robot. Similarly, $\gamma$ is calculated as the angle between the tangent of the fitted circle and the pose of the robot. Both $\theta$ and $\gamma$ are divided by the number of cycles (in our case, 3) to obtain the average angular displacement per cycle for each trial.
IV Experimental Results and Discussion
Figure 5: Experimental and simulation results of (A) heading angle per cycle and (B) curvature per cycle for a range of amplitudes ($A_{3}$). The points are the average of 5 trials and error bars are 1 standard deviation. (C) By modulating the amplitude of the third wave, different steering paths can be induced on a multi-legged robot.
IV-A Amplitude Modulation
We varied the amplitude of $A_{3}$ and plot the mean and standard deviation in Figure 5. The heading angle change and curvature ($\frac{1}{r}$) at amplitudes $A\in[0-20^{\circ}]$ demonstrate a proportionally increasing relationship with $A_{3}$ (Fig 5). The resulting average heading angle change shows an increasing trend that appears to be linearly related to $A_{3}$ until it begins to saturate after $10^{\circ}$ where the slope decreases (Fig. 5A). Additionally, due to self collisions at large body joint angles, amplitudes greater than $20^{\circ}$ are not possible yielding a maximum heading angle change of $20\pm 1.4$ degrees per cycle. Similarly, the curvature of the steering arcs generally increases with amplitude (Fig.5B), demonstrating that smaller steering radii are achievable with an inversely proportional relationship to amplitude. We observed the greatest orientation change with $A_{3}=20\textdegree$ resulting in the robot sweeping 60° over the course of 3 cycles. Interestingly, numerical simulations (Fig.5) for curvature based on GM closely match the linear relationship and bounds for the curvature while the predictions for heading change underestimate a significant portion of the graph. This is potentially due to the discrepancy between the actual ground reaction force between the robot and the carpet and to the assumed Coulomb friction model.
The arc length of the robot path is obtained by multiplying the $\theta$ by steering radius and remains relatively constant until $A_{3}=10^{\circ}$, where the total path travelled decays as total translation decreases for larger angle rotations with small turning radii. This indicates that the robot is able to maintain its translational displacement equivalent to purely forward motion ($0^{\circ}$) for small amplitude turns while its net displacement slowly decays on larger amplitude turns. The rotation angle $\gamma$ remained constant and close to zero ($<4^{\circ}$) throughout the cycle for each amplitude. While only left steering data was tracked and reported, steering right is achieved by negating $A_{3}$ and steering backwards requires changing the sign of the first temporal basis function $W_{1}$.
IV-B Spatial Basis Manipulation
Figure 6: Experimental results for (A) heading and (B) curvature for three different basis functions $S_{3}^{2}$, $S_{3}^{2}$ and $S_{3}^{3}$.(C)Snapshots from amplitude $10^{\circ}$ for $S_{3}^{2}$(C.i) and $S_{3}^{3}$(C.ii)
[…]
IV-C Obstacle Avoidance Demonstration
Figure 7: (A) Robot centipede performing an ’S’ shaped trajectory to navigate around a wall with time stamps. (B) Scolopendra polymorpha captured on a high speed camera for a low amplitude turn and (C) high amplitude turn
To validate the turning scheme as a feasible form of navigation for obstacle avoidance, we conducted a series of tests navigating around obstacles (Fig. 7.A). The demonstration of a centipede robot turning around a corner, a case where both steering right and steering left are required, shows the practical functionality of our scheme. Achieving this relatively complex path of sequencing arcs only involves changing the value of $A_{3}$ from negative to positive in between cycles while modulating the absolute value dictates the radius of the turn. This proves the effectiveness of the dimensionality reduction we propose for a high-degree-of-freedom robot control to achieve results in cluttered environments where obstacles must be avoided instead of traversed.
IV-D Turning in Multi-Legged Elongate Organisms
In addition to robot experiments, preliminary biological data was captured for a Scolopendra polymorpha on a high speed camera performing a wide variety of differing amplitude turns (Fig. 7.B, Fig. 7.C). Preliminary kinematic analysis indicates a qualitative agreement with our steering framework, but more systematic analysis and tracking is needed for validation. Using these videos, we have the initial steps towards developing a theory on how biological systems modulate their steering to implement in our robot control framework.
V Conclusion
Through the tools of geometric mechanics and previously established control techniques for robot centipedes, we effectively developed and implemented planar turning for these multi-legged robots. With the ability to control vertical and horizontal body undulation, leg rotation, and leg stepping, multi-legged robots can achieve full motion in SE(2) unlike wheeled vehicles, which are limited to the direction of their wheels. We present a novel turning scheme that supplements previous research in coordinating multi-legged robots by the addition of a second travelling wave down the body of the robot. Using simulation and robophysical experiments, we discovered that amplitude modulation of the two wave turning strategy enables the robot to effectively steer with predictable trajectories. We found two key variables in controlling steering angles: $A_{3}$, the amplitude of the second wave, and $S_{3}$, the spatial basis function. Approximating various basis functions using a fixed scaling vector allows for a smaller radius, but a theoretical framework for varying $S_{3}$ is an avenue for future work and experiments.
This scheme builds upon previous work done in centralized control of multi-legged robots involving the propagation of a single traveling wave down the body that has proven to be robust across difficult-to-traverse terrain. While rugged ground navigation using steering gaits was not investigated, preliminary outdoor tests indicated the proposed methodology to be effective. Furthermore, future research should investigate the role friction plays on turning radius predictions as these experiments were all performed in high friction carpeted environments that mitigate inertial effects. This framework creates the foundation of developing the controls that can be explored for more challenging terrain using horizontal and vertical waves to overcome obstacles.
The close agreement between the simulation results based on geometric mechanics and experiments demonstrate the ability of the proposed control mechanism to predict locomotion for multi-legged robots. This consistency highlights the predictive capabilities for moving along desired paths. As we improve robot design, particularly in the addition of segments and mechanical design, the performance of the robot is expected to increase significantly.
The predicted turning arcs serve as fundamental motion primitives that are essential for a broad range of locomotion tasks, particularly those involving autonomous navigation in complex environments and obstacle avoidance. Because of the simple amplitude modulation scheme, implementation into closed-loop control involves the manipulation of only one variable for future real-time course correction. This ability to control planar trajectories serves as a substantial step forward in the development of practical, multi-legged robot systems for real world applications.
ACKNOWLEDGMENT
We would like to thank Christopher Pierce for his technical consultation, Erik Teder for his consultation on figures and helpful discussion, and Juntao He for discussion. We would like to think the Institute for Robotics and Intelligent Machines for use of the College of Computing basement as a testing space. The authors received funding from NSF-Simons Southeast Center for Mathematics and Biology (Simons Foundation SFARI 594594), Army Research Office grant W911NF-11-1-0514, a Dunn Family Professorship, Ground Control Robotics, and a STTR Phase I (2335553) NSF grant.